# Nonreciprocal tripartite entanglement and asymmetric Einstein-Podolsky-Rosen steering via directional quantum squeezing

### v1

@@ Appendix C Derivation of the dynamics of the quantum correlation function
$M=$
($\begin{smallmatrix}-\kappa&0&0&0&0&0&0&0&0&0&0&0&-i\Lambda_{1}^{\ast}&i\Lambda_{1}&0&0\\ -i\Lambda_{2}^{\ast}&i\Lambda_{2}&0&0&0&0&0&0\\ \\ 0&-\kappa&0&0&0&0&0&0&0&0&0&0&-i\Lambda_{1}^{\ast}&i\Lambda_{1}&0&0\\ -i\Lambda_{2}^{\ast}&i\Lambda_{2}&0&0&0&0&0&0\\ \\ 0&0&0&0&0&0&0&0&2\omega_{m,1}&2\omega_{m,1}&0&0&0&0&0&0\\ 0&0&0&0&0&0&0&0\\ \\ 0&0&0&-2\gamma_{m,1}&0&0&0&0&-2\omega_{m,1}&-2\omega_{m,1}&0&0&0&0&2\Lambda_{1}^{\ast}&2\Lambda_{1}\\ 0&0&0&0&0&0&0&-2\chi\\ \\ 0&0&0&0&0&0&0&0&0&0&2\omega_{m,2}&2\omega_{m,2}&0&0&0&0\\ 0&0&0&0&0&0&0&0\\ \\ 0&0&0&0&0&-2\gamma_{m,2}&0&0&0&0&-2\omega_{m,2}&-2\omega_{m,2}&0&0&0&0\\ 0&0&2\Lambda_{2}^{\ast}&2\Lambda_{2}&0&0&-2\chi&0\\ \\ 0&0&0&0&0&0&-2\Omega_{1,+}&0&0&0&0&0&2i\Lambda_{1}&0&0&0\\ 2i\Lambda_{2}&0&0&0&0&0&0&0\\ \\ 0&0&0&0&0&0&0&2\Omega_{1,-}&0&0&0&0&0&-2i\Lambda_{1}^{\ast}&0&0\\ 0&-2i\Lambda_{2}^{\ast}&0&0&0&0&0&0\\ \\ 0&0&-\omega_{m,1}&\omega_{m,1}&0&0&0&0&-\gamma_{m,1}/2&-\gamma_{m,1}/2&0&0&\Lambda_{1}^{\ast}&\Lambda_{1}&0&0\\ 0&0&0&0&-\chi&0&0&0\\ \\ 0&0&-\omega_{m,1}&\omega_{m,1}&0&0&0&0&-\gamma_{m,1}/2&-\gamma_{m,1}/2&0&0&\Lambda_{1}^{\ast}&\Lambda_{1}&0&0\\ 0&0&0&0&-\chi&0&0&0\\ \\ 0&0&0&0&-\omega_{m,2}&\omega_{m,2}&0&0&0&0&-\gamma_{m,2}/2&-\gamma_{m,2}/2&0&0&0&0\\ \Lambda_{2}^{\ast}&\Lambda_{2}&0&0&-\chi&0&0&0\\ \\ 0&0&0&0&-\omega_{m,2}&\omega_{m,2}&0&0&0&0&-\gamma_{m,2}/2&-\gamma_{m,2}/2&0&0&0&0\\ \Lambda_{2}^{\ast}&\Lambda_{2}&0&0&-\chi&0&0&0\\ \\ 0&0&i\Lambda_{1}&0&0&0&0&0&0&0&0&0&-\Omega_{1,+}&0&\omega_{m,1}&0\\ 0&0&0&0&i\Lambda_{2}&0&0&0\\ \\ 0&0&-i\Lambda_{1}^{\ast}&0&0&0&0&0&0&0&0&0&0&\Omega_{1,-}&0&\omega_{m,1}\\ 0&0&0&0&-i\Lambda_{2}^{\ast}&0&0&0\\ \\ 0&0&0&0&0&0&0&0&0&i\Lambda_{1}&0&0&-\omega_{m,1}&0&-\Omega_{2,+}&0\\ -\chi&0&0&0&0&0&0&i\Lambda_{2}\\ \\ 0&0&0&0&0&0&0&0&0&-i\Lambda_{1}^{\ast}&0&0&0&-\omega_{m,1}&0&\Omega_{2,-}\\ 0&-\chi&0&0&0&0&0&-i\Lambda_{2}^{\ast}\\ \\ 0&0&0&0&i\Lambda_{2}&0&0&0&0&0&0&0&0&0&0&0\\ -\Omega_{1,+}&0&\omega_{m,1}&0&i\Lambda_{1}&0&0&0\\ \\ 0&0&0&0&-i\Lambda_{2}^{\ast}&0&0&0&0&0&0&0&0&0&0&0\\ 0&\Omega_{1,-}&0&\omega_{m,1}&-i\Lambda_{1}^{\ast}&0&0&0\\ \\ 0&0&0&0&0&0&0&0&0&0&0&i\Lambda_{2}&-\chi&0&0&0\\ -\omega_{m,2}&0&-\Omega_{3,+}&0&0&0&i\Lambda_{1}&0\\ \\ 0&0&0&0&0&0&0&0&0&0&0&-i\Lambda_{2}^{\ast}&0&-\chi&0&0\\ 0&-\omega_{m,2}&0&\Omega_{3,-}&0&0&-i\Lambda_{1}^{\ast}&0\\ \\ 0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0\\ 0&0&0&0&0&0&\omega_{m,2}&\omega_{m,1}\\ 0&0&0&0&0&0&0&0&0&-\chi&-\chi&0&0&0&\Lambda_{2}^{\ast}&\Lambda_{2}\\ 0&0&\Lambda_{1}^{\ast}&\Lambda_{1}&0&-\gamma_{m,1}-\gamma_{m,2}&-\omega_{m,1}&-\omega_{m,2}\\ 0&0&-\chi&0&0&0&0&0&0&0&0&0&\Lambda_{2}^{\ast}&\Lambda_{2}&0&0\\ 0&0&0&0&-\omega_{m,2}&\omega_{m,1}&-\gamma_{m,2}&0\\ 0&0&0&0&-\chi&0&0&0&0&0&0&0&0&0&0&0\\ \Lambda_{1}^{\ast}&\Lambda_{1}&0&0&-\omega_{m,1}&\omega_{m,2}&0&-\gamma_{m,1}\\ \end{smallmatrix}\right).$

### v2

@@ Appendix C Derivation of the dynamics of the quantum correlation function
$A_{11}=\left(\begin{matrix}-\kappa&0&0&0&0&0\\ 0&-\kappa&0&0&0&0\\ 0&0&0&0&0&0\\ 0&0&0&-2\gamma_{m,1}&0&0\\ 0&0&0&0&0&0\\ 0&0&0&0&0&-2\gamma_{m,2}\end{matrix}\right),$
$A_{12}=\left(\begin{matrix}0&0&0&0&0&0\\ 0&0&0&0&0&0\\ 0&0&\omega_{m,1}&\omega_{m,1}&0&0\\ 0&0&-\omega_{m,1}&-\omega_{m,1}&0&0\\ 0&0&0&0&\omega_{m,2}&\omega_{m,2}\\ 0&0&0&0&-\omega_{m,2}&-\omega_{m,2}\end{matrix}\right),$
$A_{13}=\left(\begin{matrix}-i\Lambda_{1}^{\ast}&i\Lambda_{1}&0&0&-i\Lambda_{2}^{\ast}&i\Lambda_{2}\\ -i\Lambda_{1}^{\ast}&i\Lambda_{1}&0&0&-i\Lambda_{2}^{\ast}&i\Lambda_{2}\\ 0&0&0&0&0&0\\ 0&0&2\Lambda_{1}^{\ast}&2\Lambda_{1}&0&0\\ 0&0&0&0&0&0\\ 0&0&0&0&0&0\end{matrix}\right),$
$A_{14}=\left(\begin{matrix}0&0&0&0&0&0\\ 0&0&0&0&0&0\\ 0&0&0&0&0&0\\ 0&0&0&0&0&-2\lambda\\ 0&0&0&0&0&0\\ 2\Lambda_{2}^{\ast}&2\Lambda_{2}&0&0&-2\lambda&0\end{matrix}\right),$
$A_{21}=\left(\begin{matrix}0&0&0&0&0&0\\ 0&0&0&0&0&0\\ 0&0&-\omega_{m,1}&\omega_{m,1}&0&0\\ 0&0&-\omega_{m,1}&\omega_{m,1}&0&0\\ 0&0&0&0&-\omega_{m,2}&\omega_{m,2}\\ 0&0&0&0&-\omega_{m,2}&\omega_{m,2}\end{matrix}\right),$
$A_{22}=\dfrac{1}{2}\left(\begin{matrix}-4\Omega_{1,+}&0&0&0&0&0\\ 0&4\Omega_{1,-}&0&0&0&0\\ 0&0&-\gamma_{m,1}&-\gamma_{m,1}&0&0\\ 0&0&-\gamma_{m,1}&-\gamma_{m,1}&0&0\\ 0&0&0&0&-\gamma_{m,2}&-\gamma_{m,2}\\ 0&0&0&0&-\gamma_{m,2}&-\gamma_{m,2}\end{matrix}\right),$
$A_{23}=\left(\begin{matrix}2i\Lambda_{1}&0&0&0&2i\Lambda_{2}&0\\ 0&-2i\Lambda_{1}^{\ast}&0&0&0&-2i\Lambda_{2}^{\ast}\\ \Lambda_{1}^{\ast}&\Lambda_{1}&0&0&0&0\\ \Lambda_{1}^{\ast}&\Lambda_{1}&0&0&0&0\\ 0&0&0&0&\Lambda_{2}^{\ast}&\Lambda_{2}\\ 0&0&0&0&\Lambda_{2}^{\ast}&\Lambda_{2}\end{matrix}\right),$
$A_{24}=\left(\begin{matrix}0&0&0&0&0&0\\ 0&0&0&0&0&0\\ 0&0&-\lambda&0&0&0\\ 0&0&-\lambda&0&0&0\\ 0&0&-\lambda&0&0&0\\ 0&0&-\lambda&0&0&0\end{matrix}\right),$
$A_{31}=\left(\begin{matrix}0&0&i\Lambda_{1}&0&0&0\\ 0&0&-i\Lambda_{1}^{\ast}&0&0&0\\ 0&\Lambda_{1}&0&0&0&0\\ \Lambda_{1}^{\ast}&0&0&0&0&0\\ 0&0&0&0&i\Lambda_{2}&0\\ 0&0&0&0&-i\Lambda_{2}^{\ast}&0\end{matrix}\right),$
$A_{32}=\left(\begin{matrix}0&0&0&0&0&0\\ 0&0&0&0&0&0\\ \Lambda_{1}^{\ast}&0&i\Lambda_{1}&0&0&0\\ 0&\Lambda_{1}&-i\Lambda_{1}^{\ast}&0&0&0\\ 0&0&0&0&0&0\\ 0&0&0&0&0&0\end{matrix}\right),$
$A_{33}=\left(\begin{matrix}-\Omega_{1,+}&0&\omega_{m,1}&0&0&0\\ 0&\Omega_{1,-}&0&\omega_{m,1}&0&0\\ -\omega_{m,1}&0&-\Omega_{2,+}&0&-\lambda&0\\ 0&-\omega_{m,1}&0&\Omega_{2,-}&0&-\lambda\\ 0&0&0&0&-\Omega_{1,+}&0\\ 0&0&0&0&0&\Omega_{1,-}\end{matrix}\right),$
$A_{34}=\left(\begin{matrix}0&0&i\Lambda_{2}&0&0&0\\ 0&0&-i\Lambda_{2}^{\ast}&0&0&0\\ 0&0&0&0&0&i\Lambda_{2}\\ 0&0&0&0&0&-i\Lambda_{2}^{\ast}\\ \omega_{m,1}&0&i\Lambda_{1}&0&0&0\\ 0&\omega_{m,1}&-i\Lambda_{1}^{\ast}&0&0&0\end{matrix}\right),$
$A_{41}=\left(\begin{matrix}0&\Lambda_{2}&0&0&0&0\\ \Lambda_{2}^{\ast}&0&0&0&0&0\\ 0&0&0&0&0&0\\ 0&0&0&0&0&0\\ 0&0&-\lambda&0&0&0\\ 0&0&0&0&-\lambda&0\end{matrix}\right),$
$A_{42}=\left(\begin{matrix}\Lambda_{2}^{\ast}&0&0&0&i\Lambda_{2}&0\\ 0&\Lambda_{2}&0&0&-i\Lambda_{2}^{\ast}&0\\ 0&0&0&0&0&0\\ 0&0&0&-\lambda&-\lambda&0\\ 0&0&0&0&0&0\\ 0&0&0&0&0&0\end{matrix}\right),$
$A_{43}=\left(\begin{matrix}-\lambda&0&0&0&-\omega_{m,2}&0\\ 0&-\lambda&0&0&0&-\omega_{m,2}\\ 0&0&0&0&0&0\\ 0&0&\Lambda_{2}^{\ast}&\Lambda_{2}&0&0\\ \Lambda_{2}^{\ast}&\Lambda_{2}&0&0&0&0\\ 0&0&0&0&\Lambda_{1}^{\ast}&\Lambda_{1}\end{matrix}\right),$
$A_{44}=\left(\begin{matrix}-\Omega_{3,+}&0&0&0&i\Lambda_{1}&0\\ 0&\Omega_{3,-}&0&0&-i\Lambda_{1}^{\ast}&0\\ 0&0&0&0&\omega_{m,2}&\omega_{m,1}\\ \Lambda_{1}^{\ast}&\Lambda_{1}&0&-\Omega_{4}&-\omega_{m,1}&-\omega_{m,2}\\ 0&0&-\omega_{m,2}&\omega_{m,1}&-\gamma_{m,2}&0\\ 0&0&-\omega_{m,1}&\omega_{m,2}&0&-\gamma_{m,1}\end{matrix}\right).$